Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 10,000 $ 15,000 $ 25,000
Estimated variable manufacturing overhead per machine-hour $ 1.40 $ 2.20
Job P Job Q
Direct materials $ 13,000 $ 8,000
Direct labor cost $ 21,000 $ 7,500
Actual machine-hours used:
Molding 1,700 800
Fabrication 600 900
Total 2,300 1,700

For questions 5-8 assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base.

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

5. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.)

total m cost _______

6. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.)

unit product cost _________

7. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establish selling prices for all of its jobs. What selling price would the company have established for Jobs P and Q? What are the selling prices for both jobs when stated on a per unit basis assuming 20 units were produced for Job P and 30 units were produced for Job Q? (Do not round intermediate calculations. Round your final answers to nearest whole dollar.)

Job P Job Q Total price for the job Selling price per unit

8. What was Sweeten Company’s cost of goods sold for March? (Do not round intermediate calculations.)

cost of goods sold ________
Answer #1

Answer

· Workings are given after asked Requirements

· [5]

Job Q

Direct Material

$                   8,000.00

Direct Labor

$                   7,500.00

Manufacturing Overheads applied

$                 13,515.00

Total manufacturing Cost

$                 29,015.00

· [6]

Unit Product Cost = $ 29015 / 30 = $ 967.17

· [7]

Job P

Job Q

Total manufacturing Cost

$             52,285.00

$                 29,015.00

Markup [80%] of cost

$             41,828.00

$                 23,212.00

Total Selling Price for Job

$             94,113.00

$                 52,227.00

Total units

20

30

Selling Price per unit

$               4,705.65

$                   1,740.90

· [8]

Cost of Goods Sold = $ 52285 + 29015 = $ 81,300

· Workings

Working

Molding

Fabrication

Total

A

Estimated machine hours

                        2,500

                            1,500

                 4,000

B

Variable overhead per machine hours

$                       1.40

$                           2.20

C = A x B

Total Variable Overhead

$               3,500.00

$                   3,300.00

$        6,800.00

D

Total Fixed Overhead

$             10,000.00

$                 15,000.00

$     25,000.00

E = C+D

Total Overheads

$     31,800.00

F = E/A

Plantwide Predtermined Overhead rate

$                7.95

Answer

Job P

Job Q

Manufacturing Overheads applied

$             18,285.00

$                 13,515.00

[Workings]

[2300 MHs x $ 7.95]

[1700 MHs x $ 7.95]

Job P

Direct Material

$             13,000.00

Direct Labor

$             21,000.00

Manufacturing Overheads applied

$             18,285.00

Total manufacturing Cost

$             52,285.00

· [4] Unit Product Cost = $ 52285 / 20 units
= $ 2,614.25